Math 042: Exam I Date: Oct 11, 2010

Name: Section:

1. Simplify: 8a^{3} − 3a^{2} − (4a^{2} − 5) Your answer: .

2. Solve for x: −3x + 5 = −5x − 8 Your answer: .

3. Solve for x: \frac{4}{3}x − 6 = \frac{3}{2}x Your answer: .

4. Simplify: \frac{(6x^{2}y/11)}{(9y^{2}/22)} Your answer: .

5. Evaluate: −a^{2} + 3b^{3} for a = 5, b = −3.

6. Simplify: \frac{(3xy)^{2}}{9x^{2}y^{3}} Your answer: .

7. Solve the inequality: −3(y − 1) < 12 Your answer: .

8. What is the slope of the line given by the equation x = −5?

9. Solve for x: ax + d = cx − b Your answer: .

10. If 30 is 120% of a number, what is the number? Your answer: .

11. What is 20% of 75? Your answer: .

12. 6 times the sum of a number and its cube can be represented by

(a) 6n + n^{2} (b) 6(n + n)^{3} (c) 6 + n + n^{3} (d) 6(n + n^{3})

13. The length of a rectangle is 5 less than twice its width. What can be used to represent its perimeter?

Your answer: .

14. An RCA employee earned $280 a week. If $52 is deducted for taxes, $35 for health benefits and $25 for pension, what percent of the salary is left?

(a) 168% (b) 112% (c) 60% (d) 40%

15. If $1300 is invested at t% and $600 is invested at (t − 3)% what can be used to represent the interest drawn in one year?

Your answer: .

6  Exam I  October 11, 2010

16. Which of the following is the graph of −3x + y = −6?

17. Find the slope of the line and is perpendicular to the line passing through (−2, 5) and (3, −3). Slope: .

October 11, 2010  Exam I  7

18. Find the equation of the line passing through (2, −3) with slope of 3/2. Equation: .

19. Solve the following system of equations.

2x − y = 1
−3x + 2y = 5

Your answer: .

20. Graph 3x + 2y < 6.